OpenStudy (anonymous):

Multiplying Radical Expressions Multiply: 5 the square root of 8 • 4 the square root of 7 16 the square root of 10 20 the square root of 15 40 the square root of 14 40

OpenStudy (campbell_st):

\[5\sqrt{8}\times4\times \sqrt{7 } = 5\times4\times \sqrt{8\times7}\]

OpenStudy (anonymous):

40 the square root of 14

OpenStudy (campbell_st):

just simplify

OpenStudy (anonymous):

thank you(:

OpenStudy (campbell_st):

hope it helps